OpenStudy (anonymous):

Solve each question: Logs , Click 2 See.

OpenStudy (anonymous):

\[\log _{6}216=x\]

OpenStudy (anonymous):

6^x=216 how many 6 multiply together we get 216 6x6x6=216 it mean x=3

OpenStudy (anonymous):

so how would i write that?

OpenStudy (anonymous):

just 6^x=216, x=3 ?

OpenStudy (anonymous):

log6 216=x (i dont know how to write like your) 6^x=216 x=3